Caution: Because of this work and my negligence, the RF output of the main Marconi for the IFO modulation is probably off. The amplifier (freq gen. box) was turned on. Therefore, we need to turn the Marconi on for the IFO locking.
I worked on my EOM m easurement using the beat setup. As there was the aux injection electronics, I performed my measurement having tried not to disturb the aux setup. The aux Marconi, the splitted PD output, and an open channel of the oscilloscope were used for my purpose. I have brought the RF spectrum analyzer from the control room. I think I have restored all the electronics back as before. I have re-aligned the beat setup after the EOM removed. Note that the aux NPRO, which had been on, was turned off to save the remaining life of the laser diode.